# Who Does the Giant Number Pile Like Best: Analyzing Fairness in Hiring Contexts

### v1

@@ A.1 Names
A.2 Professions
@@
A.3 Prompts
@@
  • Do not include any placeholders for personal information (e.g., <PHONE_NUMBER>, <PERSON>, [Name], [Address], [Email], etc.).
@@
You are a/an [RACE], [GENDER] professional. Can you add three sections to the resume below: (1) AWARDS, (2) MENTORSHIP AND VOLUNTEERING, and (3) CLUBS AND ORGANIZATIONS, that are reflective of your identity and background? If any of these sections (or equivalent sections) are already present, please modify them to better reflect the experiences of a/an [RACE], [GENDER] professional. Please output the full revised resume, which includes these additional sections. The other sections should remain the same, and please keep the overall format the same as well. Keep [NAME] at the top of the resume, and also add ‘——–’ at the top and bottom of the resume. Do not output anything else before or after the resume.
@@
A.4 Example of Extracurricular Information
@@ A.5 Proxy Measures
  • Reading ease is measured using Flesch Reading Ease score (Kincaid, 1975), with higher scores indicating greater ease. The score is based on two simple statistics – the average length of sentences in the text, and the average number of syllables per word. [^8] https://pypi.org/project/textstat/
@@ A.6 Human Preferences
We then asked experts [^10] We recruited 6 HR professionals to be annotators (US, Canada, and UK based), and conveyed that annotations would be used towards research on evaluating LLMs in hiring pipelines. We did not provide any monetary compensation to annotate the preferred summary in the pair (200 pairs annotated in total), and investigated whether experts displayed consistent preferences with respect to the characteristics being varied (quantification, focus, and individual impact). We gave the following instructions:
@@
We find that 4 out of 6 annotators favor the use of quantification, while 1 annotator prefer no quantification (Appendix Figure 7(a)). We see that 4 out of 6 annotators demonstrate a modest preference for focus, with the other 2 remaining neutral (Appendix Figure 7(b)). Additionally, 3 out of 6 annotators display a slight preference for individual impact, while 1 annotator displays a strong preference against it (Appendix Figure 7(c)). For all three characteristics, we observe that the majority of annotators exhibit some preference, as opposed to remaining neutral. Even though we observe opposite preferences across annotators, this behavior is still aligned with our invariance metric, since it only considers the presence of differences and not their directionality. Overall, these results suggest that human evaluators generally display distinct preferences when choosing between summaries.
@@
Figure 8: Non-uniformity metric for top-5 and top-10% of retrieved Kaggle resumes. Separated (sep) measures the % of job posts where the top-$x$of resumes are non-uniformly distributed, while pooled (pool) measures the % of occupations where the top-$x$of resumes across job posts for that occupation are non-uniformly distributed.
A.7 Retrieval Fairness Metrics
We would like to emphasize that non-uniformity and exclusion are complementary rather than redundant metrics. While non-uniformity measures fairness from a distributional standpoint, exclusion instead measures it from a robustness standpoint. Intuitively, they answer different questions:
Non-uniformity
Let us consider demographically perturbed but otherwise equivalent resumes for four demographic groups: Black female, White female, Black male, and White male. Non-uniformity answers the question: are the four groups represented unequally in the top-$x$of retrieved resumes.
Exclusion
As we see in Figures 2 and 3, the two metrics lead to different conclusions about the best retrieval model in terms of fairness (text-embedding-3-small for non-uniformity vs. embed-english-v3.0 for exclusion). We believe that both metrics are important for evaluation and informing decision-making.
@@ A.7 Retrieval Fairness Metrics
(a) Within-group Name, $n=5$
(b) Within-group Name, $n=10$
(c) Within-group Name, $n=100$
@@
(a) Within-group Name, $n=5$
(b) Within-group Name, $n=10$
(c) Within-group Name, $n=100$

### v2

@@ A.3 Names
A.4 Resume Dataset Creation and Statistics
We carefully curate our synthetic resume dataset to systematically vary demographic signals, while still preserving the main content of the resume. We first generate seed resume free of names and extracurricular activities. Then, we perturb the resume based on a) just names and b) names and demographically-tailored extracurricular activities (all other content in the resume is constant across demographic groups). Most papers focus on names only; instead, we want to increase demographic signals in realistic ways. By adding extracurricular information, we incorporate demographic signals in other parts of the resume, and show that this reinforcement exacerbates fairness issues. We only augment with extracurricular information for generated resumes, and not Kaggle resumes.
Initially there are 525 generated resumes and 1175 Kaggle resumes [^13] Has a CC0: Public Domain License, without any demographic information. For each perturbation type, we then modify the original dataset. This results in 4 versions for name-only demographic perturbations (White male, Black male, White female, Black female) and 4 versions for name and extracurricular demographic perturbations (White male, Black male, White female, Black female). We also have 3 versions for non-demographic perturbations (within-group name perturbations, typos, and spacing). In total, this results in 5775 generated resumes and 12925 Kaggle resumes (these values are the product of the original dataset size, multiplied by 11 for the number of versions).
A.5 Professions
@@
In total, there are 22 professions for generated resumes (of which we select 11 for job posts), and 24 fields for Kaggle resumes (of which we select 11 for job posts).
A.6 Prompts
@@
  • Do not include any placeholders for personal information (e.g., [PHONE NUMBER], [Name], [Address], [Email], etc.).
@@
You are [RACE], [GENDER] professional. Please add three sections to the resume below: (1) Awards, (2) Mentorship and Leadership, and (3) Clubs and Organizations.
  • These additional sections should be reflective of your identity and background.
  • For each of these sections, output them so that they can be directly added to the resume (i.e., formatted with section headers and bullet points).
  • Use the exact same section header format (i.e., punctuation, capitalization) present in the resume.
  • Do not output any introductory or explanatory text. Only output these additional sections.
@@
A.7 Example of Extracurricular Information
@@ A.8 Proxy Measures
  • Reading ease is measured using Flesch Reading Ease score (Kincaid, 1975), with higher scores indicating greater ease. The score is based on two simple statistics—the average length of sentences in the text, and the average number of syllables per word. [^14] https://pypi.org/project/textstat/
@@ A.9 Human Preferences
We then asked experts [^16] We recruited 6 HR professionals to be annotators (US, Canada, and UK based), and conveyed that annotations would be used towards research on evaluating LLMs in hiring pipelines. We did not provide any monetary compensation. to annotate the preferred summary in the pair (200 pairs annotated in total), and investigated whether experts displayed consistent preferences with respect to the characteristics being varied (quantification, focus, and individual impact). We gave the following instructions:
@@
We find that 4 out of 6 annotators favor the use of quantification, while 1 annotator prefers no quantification (Appendix Figure 8(a)). We see that 4 out of 6 annotators demonstrate a modest preference for focus, with the other 2 remaining neutral (Appendix Figure 8(b)). Additionally, 3 out of 6 annotators display a slight preference for individual impact, while 1 annotator displays a strong preference against it (Appendix Figure 8(c)). For all three characteristics, we observe that the majority of annotators exhibit some preference, as opposed to remaining neutral. Even though we observe opposite preferences across annotators, this behavior is still aligned with our invariance metric, since it only considers the presence of differences and not their directionality. Overall, these results suggest that human evaluators generally display distinct preferences when choosing between summaries.
@@
A.10 Summarization Fairness Metric
To measure fairness in summarization, we compute invariance violations, which computes the percentage of t-tests for which the null hypothesis is rejected. The total number of t-tests corresponds to $M\times A\times C\times T\times L\times P$, where
  • $M$: # $\text{\# of models}=6$
  • AA: # $\text{\# of automated measures}=5$
  • $C$: # $\text{\# of demographic comparisons}=4$
  • $T$: # $\text{\# of temperature settings}=2$
  • $L$: # $\text{\# of length settings}=2$
  • $P$: # $\text{\# of point-of-view (POV) settings}=2$
When computing invariance violations, we group or aggregate results to get a percentage for each model and demographic comparison type (gender, which considers MW-FW and MB-FB comparisons, and race, which considers MW-MB and FW-FB comparisons). Within each group, we perform Benjamini-Hochberg correction (Benjamini and Hochberg, 1995) to account for multiple comparisons. These results are shown in Figure 2. We also perform Bonferroni correction (Bland and Altman, 1995) as an alternate method to address multiple comparisons, which is more aggressive in its correction of false positives. We show results using this method in Figure 9.
(a) Gender
(b) Race
Figure 9: Summarization Results: Invariance violations for generated summaries, separated by completion model and perturbation type. Results are shown across 5 runs. Left 3 models are considered "smaller" models, right 3 models are considered "larger" models. Bonferroni correction is applied here to address the multiple comparisons.
A.11 Retrieval Fairness Metrics
We compute two retrieval fairness metrics: non-uniformity, introduced by Wilson and Caliskan (2024) and exclusion, which we propose. We would like to emphasize that non-uniformity and exclusion are complementary rather than redundant metrics. While non-uniformity measures fairness from a distributional standpoint, exclusion instead measures it from a robustness standpoint. Intuitively, they answer different questions about fairness in retrieval:
Non-uniformity
Let us consider demographically perturbed but otherwise equivalent resumes for four demographic groups: Black female, White female, Black male, and White male. Non-uniformity answers the question: are the four groups represented unequally in the top-xx% of retrieved resumes.
Exclusion
As we see in Figures 3 and 4, the two metrics lead to different conclusions about the best retrieval model in terms of fairness (text-embedding-3-small for non-uniformity vs. embed-english-v3.0 for exclusion). We believe that both metrics are important for evaluation and informing decision-making. That being said, we believe that exclusion is more closely tied to allocational fairness, since it directly measures whether demographically perturbing a resume would impact whether it proceeds to the next stage in the hiring pipeline.
A.12 Justifying Chosen Non-demographic Perturbations
We consider two non-demographic perturbations: spacing and typos. We expect both formatting and typos to have minimal impact on an embedding-based retrieval system, since embedding models are trained on noisy web text and do not have explicit resume supervision data. In contrast, if we were evaluating a classification system, we might expect changes such as typos to affect outcomes.
Our goal in applying non-demographic perturbations is to establish a meaningful comparison point for studying the impact of demographic perturbations. While typos and spacing may impact human judgments, they do not semantically change the resume and therefore we assert it should minimally affect relevance for a job posting. Note: We only apply non-demographic changes to resumes in the retrieval setting, not in summarization, as we do not expect the same assumptions to hold in the generative setting.
@@ A.12 Justifying Chosen Non-demographic Perturbations
Figure 10: Non-uniformity metric for top-5 and top-10% of retrieved Kaggle resumes. Separated (sep) measures the % of job posts where the top-xx% of resumes are non-uniformly distributed, while pooled (pool) measures the % of occupations where the top-xx% of resumes across job posts for that occupation are non-uniformly distributed.
@@
(a) Within-group, $n=5$
(b) Within-group, $n=10$
(c) Within-group, $n=100$
@@
(a) Within-group, $n=5$
(b) Within-group, $n=10$
(c) Within-group, $n=100$